OpenStudy (anonymous):

A motorboat takes 5 hours to travel 450km going upstream. The return trip takes 3 hours going downstream. What is the rate of the boat in still water and what is the rate of the current?

OpenStudy (anonymous):

The rate of the boat will be 120km/hr and the stream will b3 30km/hr. Try it.

OpenStudy (callisto):

Let rate of boat in still water be x and rate of current be y. Using the formula \(\large speed = \frac{distance}{time}\) For upstream: \[\frac{450}{x-y} = 5 \]\[450 = 5(x-y) \]\[450 = 5x-5y \ \ -(1)\] For downstream: \[\frac{450}{x+y} = 3 \]\[450 = 3(x+y)\]\[450 = 3x+3y \ \ -(2)\] (1)x3 + (2)x5 450x3 = 3(5x-5y) + 450x5 = 5(3x+3y) ------------------ 3600 = 30x Divide both sides by coefficient of x, that is 30 3600 / 30 = 30x / 30 x = 120 km/h Put x =120 into (1) 450 = 5(120) - 5y 450 = 600 - 5y -150 = -5y y = 30 km/h So, rate of boat in still water is 120km/h and rate of current is 30km/h.